# Low energy excitations in A-site ordered SmBaMn2O6

### v1

@@ Acknowledgements.
- van den Brink et al. [1999] J. van den Brink, G. Khaliullin, and D. Khomskii, Physical Review Letters 83, 5118 (1999).
- Saitoh et al. [2001] E. Saitoh, S. Okamoto, K. T. Takahashi, K. Tobe, K. Yamamoto, T. Kimura, S. Ishihara, S. Maekawa, and Y. Tokura, Nature 410, 180 (2001).
- Polli et al. [2007] D. Polli, M. Rini, S. Wall, R. W. Schoenlein, Y. Tomioka, Y. Tokura, G. Cerullo, and A. Cavalleri, Nature Materials 6, 643 (2007).
- Grüninger et al. [2002] M. Grüninger, R. Rückamp, M. Windt, P. Reutler, C. Zobel, T. Lorenz, A. Freimuth, and A. Revcolevschi, Nature 418, 39 (2002).
- Murakami et al. [1998] Y. Murakami, H. Kawada, H. Kawata, M. Tanaka, T. Arima, Y. Moritomo, and Y. Tokura, Physical Review Letters 80, 1932 (1998).
- Wilkins et al. [2003] S. B. Wilkins, P. D. Spencer, P. D. Hatton, S. P. Collins, M. D. Roper, D. Prabhakaran, and A. T. Boothroyd, Physical Review Letters 91, 167205 (2003).
- Thomas et al. [2004] K. J. Thomas, J. P. Hill, S. Grenier, Y.-J. Kim, P. Abbamonte, L. Venema, A. Rusydi, Y. Tomioka, Y. Tokura, D. F. McMorrow, G. Sawatzky, and M. van Veenendaal, Physical Review Letters 92, 237204 (2004).
- Dhesi et al. [2004] S. S. Dhesi, A. Mirone, C. De Nadaï, P. Ohresser, P. Bencok, N. B. Brookes, P. Reutler, A. Revcolevschi, A. Tagliaferri, O. Toulemonde, and G. van der Laan, Physical Review Letters 92, 056403 (2004).
- Staub et al. [2005] U. Staub, V. Scagnoli, A. M. Mulders, K. Katsumata, Z. Honda, H. Grimmer, M. Horisberger, and J. M. Tonnerre, Physical Review B 71, 214421 (2005).
- García-Fernández et al. [2008] M. García-Fernández, U. Staub, Y. Bodenthin, S. M. Lawrence, A. M. Mulders, C. E. Buckley, S. Weyeneth, E. Pomjakushina, and K. Conder, Phys. Rev. B 77, 060402 (2008).
- García-Fernández et al. [2009] M. García-Fernández, U. Staub, Y. Bodenthin, V. Scagnoli, V. Pomjakushin, S. W. Lovesey, A. Mirone, J. Herrero-Martín, C. Piamonteze, and E. Pomjakushina, Phys. Rev. Lett. 103, 097205 (2009).
- Beale et al. [2009] T. A. W. Beale, S. R. Bland, R. D. Johnson, P. D. Hatton, J. C. Cezar, S. S. Dhesi, M. v. Zimmermann, D. Prabhakaran, and A. T. Boothroyd, Physical Review B 79, 054433 (2009).
- Staub et al. [2009] U. Staub, M. García-Fernández, Y. Bodenthin, V. Scagnoli, R. A. De Souza, M. Garganourakis, E. Pomjakushina, and K. Conder, Physical Review B 79, 224419 (2009).
- García-Fernández et al. [2010] M. García-Fernández, U. Staub, Y. Bodenthin, V. Pomjakushin, A. Mirone, J. Fernández-Rodríguez, V. Scagnoli, A. M. Mulders, S. M. Lawrence, and E. Pomjakushina, Phys. Rev. B 82, 235108 (2010).
- Ishihara and Maekawa [2000] S. Ishihara and S. Maekawa, Phys. Rev. B 62, 2338 (2000).
- Ishihara et al. [2004] S. Ishihara, H. Kondoh, and S. Maekawa, Physica B: Condensed Matter 345, 15 (2004), proceedings of the Conference on Polarised Neutron and Synchrotron X-rays for Magnetism.
- Forte et al. [2008] F. Forte, L. J. P. Ament, and J. van den Brink, Phys. Rev. Lett. 101, 106406 (2008).
- Schlappa et al. [2012] J. Schlappa, K. Wohlfeld, K. J. Zhou, M. Mourigal, M. W. Haverkort, V. N. Strocov, L. Hozoi, C. Monney, S. Nishimoto, S. Singh, A. Revcolevschi, J. S. Caux, L. Patthey, H. M. Rønnow, J. van den Brink, and T. Schmitt, Nature 485, 82 (2012).
- Kondo et al. [2001] H. Kondo, S. Ishihara, and S. Maekawa, Phys. Rev. B 64, 014414 (2001).
- Inami et al. [2003] T. Inami, T. Fukuda, J. Mizuki, S. Ishihara, H. Kondo, H. Nakao, T. Matsumura, K. Hirota, Y. Murakami, S. Maekawa, and Y. Endoh, Phys. Rev. B 67, 045108 (2003).
- Ishii et al. [2004] K. Ishii, T. Inami, K. Ohwada, K. Kuzushita, J. Mizuki, Y. Murakami, S. Ishihara, Y. Endoh, S. Maekawa, K. Hirota, and Y. Moritomo, Phys. Rev. B 70, 224437 (2004).
- Grenier et al. [2005] S. Grenier, J. P. Hill, V. Kiryukhin, W. Ku, Y.-J. Kim, K. J. Thomas, S.-W. Cheong, Y. Tokura, Y. Tomioka, D. Casa, and T. Gog, Phys. Rev. Lett. 94, 047203 (2005).
- Nakajima et al. [2004a] T. Nakajima, H. Yoshizawa, and Y. Ueda, Journal of the Physical Society of Japan 73, 2283 (2004a), https://doi.org/10.1143/JPSJ.73.2283 .
- Uchida et al. [2002] M. Uchida, D. Akahoshi, R. Kumai, Y. Tomioka, T.-h. Arima, Y. Tokura, and Y. Matsui, Journal of the Physical Society of Japan 71, 2605 (2002), https://doi.org/10.1143/JPSJ.71.2605 .
- Kageyama et al. [2003] H. Kageyama, T. Nakajima, M. Ichihara, Y. Ueda, H. Yoshizawa, and K. Ohoyama, Journal of the Physical Society of Japan 72, 241 (2003), https://doi.org/10.1143/JPSJ.72.241 .
- Nakajima et al. [2004b] T. Nakajima, H. Kageyama, M. Ichihara, K. Ohoyama, H. Yoshizawa, and Y. Ueda, Journal of Solid State Chemistry 177, 987 (2004b).
- Arima et al. [2002] T. Arima, D. Akahoshi, K. Oikawa, T. Kamiyama, M. Uchida, Y. Matsui, and Y. Tokura, Phys. Rev. B 66, 140408 (2002).
- Akahoshi et al. [2004] D. Akahoshi, Y. Okimoto, M. Kubota, R. Kumai, T. Arima, Y. Tomioka, and Y. Tokura, Phys. Rev. B 70, 064418 (2004).
- Morikawa et al. [2012] D. Morikawa, K. Tsuda, Y. Maeda, S. Yamada, and T.-h. Arima, Journal of the Physical Society of Japan 81, 093602 (2012), https://doi.org/10.1143/JPSJ.81.093602 .
- Tokunaga et al. [2006] Y. Tokunaga, T. Lottermoser, Y. Lee, R. Kumai, M. Uchida, T. Arima, and Y. Tokura, Nature Materials 5, 937 (2006).
- Zhou et al. [2022] K.-J. Zhou, A. Walters, M. Garcia-Fernandez, T. Rice, M. Hand, A. Nag, J. Li, S. Agrestini, P. Garland, H. Wang, S. Alcock, I. Nistea, B. Nutter, N. Rubies, G. Knap, M. Gaughran, F. Yuan, P. Chang, J. Emmins, and G. Howell, Journal of Synchrotron Radiation 29, 563 (2022).
- Ewings et al. [2016] R. A. Ewings, T. G. Perring, O. Sikora, D. L. Abernathy, Y. Tomioka, and Y. Tokura, Phys. Rev. B 94, 014405 (2016).
- Ishihara et al. [2005] S. Ishihara, Y. Murakami, T. Inami, K. Ishii, J. Mizuki, K. Hirota, S. Maekawa, and Y. Endoh, New Journal of Physics 7, 119 (2005).
- van den Brink [2001] J. van den Brink, Physical Review Letters 87, 217202 (2001).
- Allen and Perebeinos [1999] P. B. Allen and V. Perebeinos, Physical Review Letters 83, 4828 (1999).
- Schmidt et al. [2007] K. P. Schmidt, M. Grüninger, and G. S. Uhrig, Physical Review B 76, 075108 (2007).

### v2

@@ Acknowledgements.
- [1] (2012) Charge and orbital order patterns in an a-site ordered perovskite-type manganite smbamn2o6 determined by convergent-beam electron diffraction. Journal of the Physical Society of Japan 81 (9), pp. 093602. Cited by: §I, §I.
- [2] (2003) New stacking variations of the charge and orbital ordering in the metal-ordered manganite ybamn2o6. Journal of the Physical Society of Japan 72 (2), pp. 241–244. Cited by: §I.
- [3] (2002) Charge/orbital ordering structure in ordered perovskite sm1/2ba1/2mno3. Journal of the Physical Society of Japan 71 (11), pp. 2605–2608. Cited by: §I.
- [4] (2004) A-site randomness effect on structural and physical properties of ba-based perovskite manganites. Journal of the Physical Society of Japan 73 (8), pp. 2283–2291. Cited by: §I.
- [5] (2004-08) Charge-orbital ordering near the multicritical point in -site ordered perovskites and . Phys. Rev. B 70, pp. 064418. Cited by: §I, §II, §III.
- [6] (1999-12) Self-trapped exciton and franck-condon spectra predicted in lamno3. Physical Review Letters 83 (23), pp. 4828–4831. Cited by: §III.
- [7] (2002-10) Change in charge and orbital alignment upon antiferromagnetic transition in the a-site-ordered perovskite manganese oxide and sm). Phys. Rev. B 66, pp. 140408. Cited by: §I.
- [8] (2009-02) Thermally induced rotation of 3d orbital stripes in prsr0.1ca0.9mno7. Physical Review B 79 (5), pp. 054433–. Cited by: §I.
- [9] (2004-02) Unraveling orbital ordering in la0.5sr1.5mno4. Physical Review Letters 92 (5), pp. 056403–. Cited by: §I.
- [10] (2016-07) Spin excitations used to probe the nature of exchange coupling in the magnetically ordered ground state of pr0.5ca0.5mno3. Phys. Rev. B 94, pp. 014405. Cited by: §III.
- [11] (2008-09) Single and double orbital excitations probed by resonant inelastic x-ray scattering. Phys. Rev. Lett. 101, pp. 106406. Cited by: §I, §III, §III.
- [12] (2008-02) Resonant soft x-ray powder diffraction study to determine the orbital ordering in a-site-ordered . Phys. Rev. B 77, pp. 060402. Cited by: §I, §I, §III, §III, §III.
- [13] (2010-12) Doping and temperature dependence of states in -site ordered manganites. Phys. Rev. B 82, pp. 235108. Cited by: §I, §III, §III.
- [14] (2009-08) Orbital order at mn and o sites and absence of zener polaron formation in manganites. Phys. Rev. Lett. 103, pp. 097205. Cited by: §I, §I, §III, §III.
- [15] (2005-02) Excitations in manganites probed by resonant inelastic x-ray scattering. Phys. Rev. Lett. 94, pp. 047203. Cited by: §I.
- [16] (2002) Experimental quest for orbital waves. Nature 418 (6893), pp. 39–40. Cited by: §I.
- [17] (2003-01) Orbital excitations in studied by resonant inelastic x-ray scattering. Phys. Rev. B 67, pp. 045108. Cited by: §I.
- [18] (2004) Resonant inelastic x-ray scattering in manganites with perovskite structure. Physica B: Condensed Matter 345 (1), pp. 15–18. Cited by: §I.
- [19] (2000-07) Theory of orbital excitation and resonant inelastic x-ray scattering in manganites. Phys. Rev. B 62, pp. 2338–2345. Cited by: §I.
- [20] (2005) Theory and experiment of orbital excitations in correlated oxides. New Journal of Physics 7 (1), pp. 119. Cited by: §III.
- [21] (2004-12) Resonant inelastic x-ray scattering study of the hole-doped manganites (, 0.4). Phys. Rev. B 70, pp. 224437. Cited by: §I.
- [22] (2001-06) Resonant inelastic x-ray scattering from charge and orbital excitations in manganites. Phys. Rev. B 64, pp. 014414. Cited by: §I.
- [23] (1998-03) Direct observation of charge and orbital ordering in la0.5sr1.5mno4. Physical Review Letters 80 (9), pp. 1932–1935. Cited by: §I.
- [24] (2004) Anomalous octahedral distortion and multiple phase transitions in the metal-ordered manganite ybamn2o6. Journal of Solid State Chemistry 177 (3), pp. 987–999. Cited by: §I.
- [25] (2007) Coherent orbital waves in the photo-induced insulator–metal dynamics of a magnetoresistive manganite. Nature Materials 6 (9), pp. 643–647. Cited by: §I.
- [26] (2001) Observation of orbital waves as elementary excitations in a solid. Nature 410 (6825), pp. 180–183. Cited by: §I.
- [27] (2012) Spin–orbital separation in the quasi-one-dimensional mott insulator sr2cuo3. Nature 485 (7396), pp. 82–85. Cited by: §I.
- [28] (2007-08) Fate of orbitons coupled to phonons. Physical Review B 76 (7), pp. 075108–. Cited by: §III, §III.
- [29] (2009-06) Orbital and magnetic ordering in and manganites near half doping studied by resonant soft x-ray powder diffraction. Physical Review B 79 (22), pp. 224419–. Cited by: §I.
- [30] (2005-06) Orbital and magnetic ordering in la0.5sr1.5mno4 studied by soft x-ray resonant scattering. Physical Review B 71 (21), pp. 214421–. Cited by: §I.
- [31] (2004-06) Soft x-ray resonant diffraction study of magnetic and orbital correlations in a manganite near half doping. Physical Review Letters 92 (23), pp. 237204–. Cited by: §I.
- [32] (2006) Rotation of orbital stripes and the consequent charge-polarized state in bilayer manganites. Nature Materials 5 (12), pp. 937–941. Cited by: §I.
- [33] (2023) Chiral phonons in quartz probed by x-rays. Nature 618 (7967), pp. 946–950. Cited by: §III.
- [34] (1999-12) Charge and orbital order in half-doped manganites. Physical Review Letters 83 (24), pp. 5118–5121. Cited by: §I.
- [35] (2001-11) Orbital excitations in lamno3. Physical Review Letters 87 (21), pp. 217202–. Cited by: §III.
- [36] (2003-10) Direct observation of orbital ordering in la0.5sr1.5mno4 using soft x-ray diffraction. Physical Review Letters 91 (16), pp. 167205–. Cited by: §I.
- [37] (2022-03) I21: an advanced high-resolution resonant inelastic x-ray scattering beamline at diamond light source. Journal of Synchrotron Radiation 29 (2), pp. 563–580. Cited by: §II.